Hi, i need help with this question!
A group consists of 5 democrats, 6 republicans, and 4 independents. Two group members are randomly selected in succession. Find the probability of selecting an Independent and then a democrat.
This is what I got so far...
# of democrats= 5, # of republicans= 6, # of independents = 4 TOTAL = 15 members
probability of selecting Independent 15C4/15
probability of selecting Democrat 15C5/15
** this is where i need help, i dont know where to go from here!

Answer by stanbon(75887)

A group consists of 5 democrats, 6 republicans, and 4 independents. Two group members are randomly selected in succession. Find the probability of selecting an Independent and then a democrat.
---
"In succession"implies without replacement
This is what I got so far...
# of democrats= 5, # of republicans= 6, # of independents = 4 TOTAL = 15 members
probability of selecting Independent [4C1]/15 = 4/15
probability of selecting Democrat after selecting an independent = 5C1/14
---------------------------------
P(inde and then democrat) = (4/15)*(5/14) = 2/21
==================================================
Cheers,
Stan H.